# Safe Control of Multi-Agent Systems with Minimal Communication

### v1

@@ I Introduction
Notation. The set of positive integers is denoted as $\mathbb{N}_{+}$. We denote by II the identity matrix, with dimension determined by context unless explicitly stated otherwise. The notation $\text{blkdiag}(A_{1}\,\ldots,\,A_{n})$ denotes a diagonal block matrix with diagonal blocks $A_{1},\,\ldots,\,A_{n}$. Given an index set $\mathcal{N}$, we write $\mathcal{N}_{-i}$ to mean the set excluding the element $i$, i.e., $\mathcal{N}\setminus\{i\}$. For a matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$, $M(i,j)$ denotes its element in the $i$th row and the $j$th column.
@@ II Problem Statement
Consider a network of NN agents indexed by the set $\mathcal{N}:=\{1,\,2,\,\ldots,\,N\}$. Each agent $i$ has decoupled linear time-varying dynamics with coupled measurements
@@
over a finite horizon $T$, where $x^{i}_{t}\in\mathbb{R}^{n_{x}^{i}}$, $u^{i}_{t}\in\mathbb{R}^{n^{i}_{u}}$, $w^{i}_{t}\in\mathbb{R}^{n^{i}_{x}}$, $y^{i}_{t}\in\mathbb{R}^{n^{i}_{y}}$ and $v^{i}_{t}\in\mathbb{R}^{n^{i}_{y}}$ represent local state vector, control action, state disturbances, measurement, and measurement noise, respectively. The state dynamics of each agent is decoupled and time-varying, driven by $A]^{ii}_{t}$ and $B]^{ii}_{t}$. The set $\mathcal{O}_{i}$ contains the indices associated with the agents that influence the measurements of agent ii.
We use $x_{t}\in\mathbb{R}^{n_{x}}$, $u_{t}\in\mathbb{R}^{n_{u}}$, $w_{t}\in\mathbb{R}^{n_{x}}$, $y_{t}\in\mathbb{R}^{n_{y}}$ and $v_{t}\in\mathbb{R}^{n_{y}}$ to denote the joint vectors of the NN agents with $x_{t}=\begin{bmatrix}\left(x_{t}^{1}\right)^{\top};\left(x_{t}^{2}\right)^{\top};\cdots;\left(x_{t}^{N}\right)^{\top}\end{bmatrix}^{\top}$. The concatenated vectors $u_{t}$, $w_{t}$, $y_{t}$, $v_{t}$ are defined similarly. We define the global dynamics matrices at time $t$, denoted by $A_{t}$, $B_{t}$, and $C_{t}$, by assembling the submatrices $A]^{ij}_{t}$, $B]^{ij}_{t}$, and $C]^{ij}_{t}$ into block matrices. Specifically, each $\cdot\,]_{t}^{ij}$ is placed in the $(i,j)$th block of the corresponding global matrix.
@@
where $e^{ij}_{(k,\tau)}\in\mathbb{R}^{n^{j}_{y}}$ serves as a linear encoder at agent $j$, and each $m^{ij}_{k}\in\mathbb{R}$ is an encoded message sent from agent $j$ to agent $i$ at time $t_{k}$. The number of messages sent from $j$ to $i$ is $r^{ij}$. The contribution of agent $j$’s information to $u^{i}_{t}$ is then computed by using a local decoder $d^{ij}_{(t,k)}\in\mathbb{R}^{n^{i}_{u}}$ at agent ii. This controller structure is illustrated in Figure 1.
@@ IV Numerical Evaluation
To evaluate the performance of the proposed minimal communication controller, we consider a multi-agent system composed of NN vehicles, each modeled as a two-dimensional double integrator (dropping the agent indexing $i$ for brevity of notation): $\ddot{p}_{x}=u_{x}$, $\ddot{p}_{y}=u_{y}$, where $(p_{x},p_{y})$ represent the $(x,y)$ position of the vehicle driven by the force $(u_{x},u_{y})$. The state of each vehicle is $x=\begin{bmatrix}p_{x}&p_{y}&\dot{p}_{x}&\dot{p}_{y}\end{bmatrix}^{\top}$ . We discretize the vehicle dynamics with unit step over a finite time horizon $T=10$. For different tasks, we shall specify different measurement neighbors $\mathcal{O}_{i}$ and measurement parameter matrices $C]^{ij}_{t}$ for all $i,\,j\in\mathcal{N}$ and $t\leq 10$. We add independently and identically distributed uniform disturbance $w_{t}$ and noise $v_{t}$ to each agent. Unless otherwise specified, we set all coordinates of the disturbances $w_{t}$ to be in range [0.05,0.05][-0.05,0.05] and $v_{t}$ to be in range [0.05,0.05][-0.05,0.05] for all agents. For all experiments, the global control actions are constrained to be within $u\in[-2,2]^{2N}$.
@@ IV-B Trajectory tracking with various sensing scenarios
The two vehicles are required to maintain an $L_{1}$ distance of $d_{t}(1,2)=15$ for decoupled measurements, 1515 for relative measurements, and 1414 for heterogeneous sensors experiments, for all $t\in{1,\ldots,10}$.
@@ V Conclusion
We now consider the term $\mathbf{\Phi}^{ux}\left(\mathbf{\Phi}^{xx}\right)^{-1}\mathbf{\Phi}^{xy}$. First, we claim that $\left(\mathbf{\Phi}^{xx}\right)^{-1}\in(\mathcal{D}+\mathcal{L})$. To see this, note that due to the affine subspace constraint (10), $\mathbf{\Phi}^{xx}$ is required to have $\Phi^{xx}_{(t,t)}=I$ for all $t\in\{0,\ldots,T\}$. Therefore, $\mathbf{\Phi}^{xx}=I+N$ with $I\in\mathbb{R}^{(T+1)n_{x}\times(T+1)n_{x}}$ and NN a block strictly lower triangular matrix, which is nilpotent. In particular, there exists $n\leq T+1$ such that $N^{n}=0$. Therefore, we can write $\left(\mathbf{\Phi}^{xx}\right)^{-1}=I+\sum_{k=1}^{n-1}(-1)^{k}N^{k}$. Since multiplication and summation of block lower triangular matrices remain block lower triangular, $\left(\mathbf{\Phi}^{xx}\right)^{-1}\in(\mathcal{D}+\mathcal{L})$. Furthermore, note that $\mathbf{\Phi}^{xx,21}$ corresponds to the entries of $\mathbf{\Phi}^{xx}$ supported on the sparsity pattern defined by $\mathcal{L}$. Referring to Table II, which can be verified via simple calculation, we see that if $A\in\mathcal{D}$ and $B\in\mathcal{L}$, then $A^{2}\in\mathcal{D}$, $B^{2}=0$, and $AB\in\mathcal{L}$. Therefore, for all $k\in\{1,\ldots,n\}$, the entries of $N^{k}$ supported on the sparsity pattern defined by $\mathcal{L}$ are all linear combination of summation of multiplication of $\mathbf{\Phi}^{xx,21}$. Consequently, the same holds for $\left(\mathbf{\Phi}^{xx}\right)^{-1}$. Using F1 and F3, the rank of $\mathcal{L}\left(\left(\mathbf{\Phi}^{xx}\right)^{-1}\right)$ is at most $\text{rank}(\mathbf{\Phi}^{xx,21})$.

### v2

@@ I Introduction
Notation. The set of positive integers is denoted as $\mathbb{N}_{+}$. We denote by $I$ the identity matrix, with dimension determined by context unless explicitly stated otherwise. The notation $\text{blkdiag}(A_{1}\,\ldots,\,A_{n})$ denotes a diagonal block matrix with diagonal blocks $A_{1},\,\ldots,\,A_{n}$. Given an index set $\mathcal{N}$, we write $\mathcal{N}_{-i}$ to mean the set excluding the element $i$, i.e., $\mathcal{N}\setminus\{i\}$. For a matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$, $M(i,j)$ denotes its element in the $i$th row and the $j$th column.
@@ II Problem Statement
Consider a network of $N$ agents indexed by the set $\mathcal{N}:=\{1,\,2,\,\ldots,\,N\}$. Each agent $i$ has decoupled linear time-varying dynamics with coupled measurements
@@
over a finite horizon $T$, where $x^{i}_{t}\in\mathbb{R}^{n_{x}^{i}}$, $u^{i}_{t}\in\mathbb{R}^{n^{i}_{u}}$, $w^{i}_{t}\in\mathbb{R}^{n^{i}_{x}}$, $y^{i}_{t}\in\mathbb{R}^{n^{i}_{y}}$ and $v^{i}_{t}\in\mathbb{R}^{n^{i}_{y}}$ represent local state vector, control action, state disturbances, measurement, and measurement noise, respectively. The state dynamics of each agent is decoupled and time-varying, driven by $A]^{ii}_{t}$ and $B]^{ii}_{t}$. The set $\mathcal{O}_{i}$ contains the indices associated with the agents that influence the measurements of agent $i$.
We use $x_{t}\in\mathbb{R}^{n_{x}}$, $u_{t}\in\mathbb{R}^{n_{u}}$, $w_{t}\in\mathbb{R}^{n_{x}}$, $y_{t}\in\mathbb{R}^{n_{y}}$ and $v_{t}\in\mathbb{R}^{n_{y}}$ to denote the joint vectors of the $N$ agents with $x_{t}=\begin{bmatrix}\left(x_{t}^{1}\right)^{\top};\left(x_{t}^{2}\right)^{\top};\cdots;\left(x_{t}^{N}\right)^{\top}\end{bmatrix}^{\top}$. The concatenated vectors $u_{t}$, $w_{t}$, $y_{t}$, $v_{t}$ are defined similarly. We define the global dynamics matrices at time $t$, denoted by $A_{t}$, $B_{t}$, and $C_{t}$, by assembling the submatrices $A]^{ij}_{t}$, $B]^{ij}_{t}$, and $C]^{ij}_{t}$ into block matrices. Specifically, each $\cdot\,]_{t}^{ij}$ is placed in the $(i,j)$th block of the corresponding global matrix.
@@
where $e^{ij}_{(k,\tau)}\in\mathbb{R}^{n^{j}_{y}}$ serves as a linear encoder at agent $j$, and each $m^{ij}_{k}\in\mathbb{R}$ is an encoded message sent from agent $j$ to agent $i$ at time $t_{k}$. The number of messages sent from $j$ to $i$ is $r^{ij}$. The contribution of agent $j$’s information to $u^{i}_{t}$ is then computed by using a local decoder $d^{ij}_{(t,k)}\in\mathbb{R}^{n^{i}_{u}}$ at agent $i$. This controller structure is illustrated in Figure 1.
@@ IV Numerical Evaluation
To evaluate the performance of the proposed minimal communication controller, we consider a multi-agent system composed of $N$ vehicles, each modeled as a two-dimensional double integrator (dropping the agent indexing $i$ for brevity of notation): $\ddot{p}_{x}=u_{x}$, $\ddot{p}_{y}=u_{y}$, where $(p_{x},p_{y})$ represent the $(x,y)$ position of the vehicle driven by the force $(u_{x},u_{y})$. The state of each vehicle is $x=\begin{bmatrix}p_{x}&p_{y}&\dot{p}_{x}&\dot{p}_{y}\end{bmatrix}^{\top}$ . We discretize the vehicle dynamics with unit step over a finite time horizon $T=10$. For different tasks, we shall specify different measurement neighbors $\mathcal{O}_{i}$ and measurement parameter matrices $C]^{ij}_{t}$ for all $i,\,j\in\mathcal{N}$ and $t\leq 10$. We add independently and identically distributed uniform disturbance $w_{t}$ and noise $v_{t}$ to each agent. Unless otherwise specified, we set all coordinates of the disturbances $w_{t}$ to be in range [0.05,0.05][-0.05,0.05] and $v_{t}$ to be in range [0.05,0.05][-0.05,0.05] for all agents. For all experiments, the global control actions are constrained to be within $u\in[-2,2]^{2N}$.
@@ IV-B Trajectory tracking with various sensing scenarios
The two vehicles are required to maintain an $L_{1}$ distance of $d_{t}(1,2)=15$ for decoupled measurements, $15$ for relative measurements, and $14$ for heterogeneous sensors experiments, for all $t\in{1,\ldots,10}$.
@@ V Conclusion
We now consider the term $\mathbf{\Phi}^{ux}\left(\mathbf{\Phi}^{xx}\right)^{-1}\mathbf{\Phi}^{xy}$. First, we claim that $\left(\mathbf{\Phi}^{xx}\right)^{-1}\in(\mathcal{D}+\mathcal{L})$. To see this, note that due to the affine subspace constraint (10), $\mathbf{\Phi}^{xx}$ is required to have $\Phi^{xx}_{(t,t)}=I$ for all $t\in\{0,\ldots,T\}$. Therefore, $\mathbf{\Phi}^{xx}=I+N$ with $I\in\mathbb{R}^{(T+1)n_{x}\times(T+1)n_{x}}$ and $N$ a block strictly lower triangular matrix, which is nilpotent. In particular, there exists $n\leq T+1$ such that $N^{n}=0$. Therefore, we can write $\left(\mathbf{\Phi}^{xx}\right)^{-1}=I+\sum_{k=1}^{n-1}(-1)^{k}N^{k}$. Since multiplication and summation of block lower triangular matrices remain block lower triangular, $\left(\mathbf{\Phi}^{xx}\right)^{-1}\in(\mathcal{D}+\mathcal{L})$. Furthermore, note that $\mathbf{\Phi}^{xx,21}$ corresponds to the entries of $\mathbf{\Phi}^{xx}$ supported on the sparsity pattern defined by $\mathcal{L}$. Referring to Table II, which can be verified via simple calculation, we see that if $A\in\mathcal{D}$ and $B\in\mathcal{L}$, then $A^{2}\in\mathcal{D}$, $B^{2}=0$, and $AB\in\mathcal{L}$. Therefore, for all $k\in\{1,\ldots,n\}$, the entries of $N^{k}$ supported on the sparsity pattern defined by $\mathcal{L}$ are all linear combination of summation of multiplication of $\mathbf{\Phi}^{xx,21}$. Consequently, the same holds for $\left(\mathbf{\Phi}^{xx}\right)^{-1}$. Using F1 and F3, the rank of $\mathcal{L}\left(\left(\mathbf{\Phi}^{xx}\right)^{-1}\right)$ is at most $\text{rank}(\mathbf{\Phi}^{xx,21})$.